How do you solve these linear inequalities algebraically?
3x+4y=5
Ax+3y=B

Found 2 solutions by Sanskriti, longjonsilver:
Answer by Sanskriti(1)

Plug in the point(-3,4) in the equation for x and y.
A(-3) + B(4) = 11
-3A + 4B = 11------(1)
Plug in the point (2,1) in the equation.
A(2) + B(1) = 11
2A + B = 11------(2)
From (2) 2A + B = 11
Subtracting 2A on both sides,
2A + B – 2A = 11 – 2A
B = 11 – 2 A
Plugging this in (1), we get,
-3A + 4( 11-2A = 11
-3A + 44 – 8A = 11
-11A = – 33
A = 3
Plug in the value of A in any one of the equation (1) and (2)
2A + B = 11
2(3) + B = 11
6 + B = 11
B = 11 – 6 = 5.
So A = 3 and B =5 is the solution.



Answer by longjonsilver(2297)

i have no idea what you are asking here..you have 4 things you do not know, x, y, A and B but only 2 equations...no way to solve this.

As for the Sanskriti solution, it is jibberish, ignore it!.

Jon.